OpenStudy (superhelp101):

Find the solution to the equation 64^(4 – x) = 4^2x. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (superhelp101):

@SolomonZelman

OpenStudy (imstuck):

Take the log of both sides, like this:

OpenStudy (imstuck):

\[\log64^{(4-x)}=\log4^{2x}\]

OpenStudy (imstuck):

Then bring down the exponents as the rules tell us:\[(4-x)\log64=(2x)\log4\]

OpenStudy (superhelp101):

oh ok

OpenStudy (imstuck):

Then...\[\frac{ 4-x }{ 2x }=\frac{ \log4 }{ \log64 }\]

OpenStudy (imstuck):

\[\frac{ 4-x }{ 2x }=.333\]\[4-x=2x(.333)\]\[4-x=.666x\]\[4=.666x+x\]\[4=1.666x\]\[2.4=x\]

OpenStudy (imstuck):

Can you translate that into complete sentences using what I showed you?

OpenStudy (superhelp101):

yep, that makes sense. Thank you for the help :)